# PASER [Uncaptioned image]: Post-Training Data Selection for Efficient Pruned Large Language Model Recovery

### v1

@@ Abstract
Model pruning is an effective approach for compressing large language models. However, this process often leads to significant degradation of model capabilities. While post-training techniques such as instruction tuning are commonly employed to recover model performance, existing methods often overlook the uneven deterioration of model capabilities and incur high computational costs. Moreover, some instruction data irrelevant to model capability recovery may introduce negative effects. To address these challenges, we propose the Post-training dAta Selection method for Efficient pruned large language model Recovery (PASER). PASER aims to identify instructions where model capabilities are most severely compromised within a certain recovery data budget. Our approach first applies manifold learning and spectral clustering to group recovery data in the semantic space, revealing capability-specific instruction sets. We then adaptively allocate the data budget to different clusters based on the degrees of model capability degradation. In each cluster, we prioritize data samples where model performance has declined dramatically. To mitigate potential negative transfer, we also detect and filter out conflicting or irrelevant recovery data. Extensive experiments demonstrate that PASER significantly outperforms conventional baselines, effectively recovering the general capabilities of pruned LLMs while utilizing merely 4%-20% of the original post-training data.
Machine Learning, ICML
@@ 1 Introduction
Model pruning, which aims at reducing model parameter amounts while maintaining model capabilities, has been a promising approach for large language model (LLM) compression. Mainstream LLM pruning schemes include unstructured (Frantar & Alistarh, 2023), semi-structured (Sun et al., 2024), and structured pruning (Ma et al., 2023). In practice, the capability degradation of the pruned model compared with the original LLM is almost unavoidable, especially under high pruning ratios. This degradation phenomenon is often more severe for the structured pruning scheme (Dong et al., 2024), which has been widely adopted in industrial LLM compression thanks to its hardware-friendly property. Therefore, first pruning, then recovery post-training has been a standard pipeline (Ma et al., 2023; Zhao et al., ). Among various types of data including pre-training corpora and extensive fine-tuning datasets (Xia et al., 2024; Sun et al., 2024), instruction tuning data has demonstrated unique advantages for efficient capability recovery (Ma et al., 2023; Zhao et al., ; Zhang et al., 2024; Chen et al., 2023). Compared to pre-training which requires massive computational resources, instruction tuning enables effective recovery with a much smaller data scale by explicit supervision. Furthermore, through the diverse task coverage, like language modeling, common sense reasoning, mathematical problem solving, and code generation, instruction tuning preserves the model’s general-purpose capabilities while preventing over-specialized recovery.
@@
Conventional schemes (Ma et al., 2023) usually employ the full version of publicly available instruction tuning datasets like Alpaca (Taori et al., 2023) to conduct the recovery post-training. However, this can bring significant computation overhead and even unsatisfactory recovery performance (See Appendix A). An intuitive solution is to take part of the original data for training, thus consuming less data and correspondingly reducing the computation resource demand. Nevertheless, directly utilizing the uniformly split data subset (e.g., first 20% of the data), can lead to sub-optimal performance, or even performance degradation. Moreover, the recovered performance considerably varies for models trained with different subsets. Therefore, selecting the most valuable instruction-tuning data that can contribute to recovery performance and reduce training costs becomes crucial. Though previous works have noticed the significance of selecting high-quality data to conduct the general instruction tuning (Wang et al., 2024), few of them are specifically designed for the recovery post-training. Note that general high quality does not necessarily mean useful for recovery.
@@
Considering the above limitations, the ideal recovery training data selection approach should exhibit the following properties: (1) Targeted and Balanced Capability Recovery: Given the uneven deterioration of different capabilities in the pruning (see Appendix A), the ideal selection method should identify and prioritize the severely-degraded ones, while ensuring balanced recovery of the model’s overall functionality. Therefore, it needs to select a comprehensive instruction set that effectively targets the most affected capabilities while still allocating appropriate recovery data to less impacted ones. (2) Recovery Training Efficiency: Limited computing resources pose serious efficiency challenges to the LLM recovery post-training. An ideal method should be able to select instructions that are both most beneficial for recovery and low in computational cost, thereby accelerating the training process and optimizing resource utilization. (3) Mitigation of Negative Transfer: Recognizing that not all instruction data is beneficial for model recovery, an optimal approach should not only identify the most advantageous instructions, but also filter out potentially harmful or irrelevant ones. This significantly reduces the risk of negative transfer during the recovery training, ensuring that the selected data contributes positively for model recovery.
@@
To achieve such goals, we propose the Post-training dAta Selection method for Efficient pruned large language model Recovery (PASER). First, we perform semantic-structural recovery instruction clustering to obtain different data groups corresponding to different LLM capabilities. Second, we select recovery instructions in a capability degradation-aware manner, where the overall data budget is allocated to different clusters based on their corresponding capability degradation degrees. In particular, when selecting samples within each capability cluster, the computation cost of each sample is also taken into consideration to optimize the efficiency of the recovery process. Finally, we construct the concept consistency graph to maintain the semantic consistency across selected instructions, thus preventing introducing conflicting or irrelevant samples. We take the LLaMA 2/3 and Baichuan2 as the target LLMs and perform the experiments under different LLM pruning schemes and different-sized instruction tuning datasets. The comparison with random and conventional instruction tuning data selection baselines demonstrates that PASER can more effectively enhance the recovered LLM performance on language modeling and various reasoning tasks. Meanwhile, the recovery training overhead can also be reduced significantly.
@@ 2 Related Works
Large Language Model Pruning can be generally divided into three categories: unstructured pruning, semi-structured pruning, and structured pruning. Unstructured pruning removes individual weights without structural constraints, with representative works including SparseGPT (Frantar & Alistarh, 2023), Wanda (Sun et al., 2024), BESA (Xu et al., 2024b), and OWL (Yin et al., 2024). This technique allows for maximum flexibility in weight selection and can achieve high compression rates while maintaining model performance. However, the resulting irregular sparsity patterns limits the practical acceleration. Semi-structured pruning (Guo et al., 2024; Malla et al., 2024; Frantar & Alistarh, 2023; Sun et al., 2024) targets specific patterns like N:M sparsity, balancing flexibility and hardware efficiency. Structured pruning approaches like LLM-Pruner (Ma et al., 2023) and SliceGPT (Ashkboos et al., 2024) remove entire structural components, offering better hardware compatibility and attracting industry attention (Ko et al., 2023; An et al., 2024; Song et al., 2024; Xia et al., 2024). However, structured pruning faces more severe performance degradation, highlighting the importance of recovery post-training.
@@ 3 Methodology
In this section, we first formulate the problem and then introduce three main components of the PASER framework (shown in Figure 1): semantic-structural recovery instruction clustering, capability degradation-aware instruction selection, and negative transfer mitigation. Furthermore, we provide the time complexity analysis for PASER process.
Figure 1: Overall Framework for our proposed PASER recovery data selection framework.
@@ 3.1 Problem Formulation
Let $M_{o}$ denote the original large language model and $M_{p}$ the pruned version of this model. We define the instruction tuning dataset as $D=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ represents an instruction and $y_{i}$ its corresponding output. Our goal is to select a subset $S\subset D$ to efficiently recover the performance of $M_{p}$. We formulate the problem as an optimization task:
@@ 3.2 Semantic-Structural Recovery Instruction Clustering
During the LLM pruning process, different model capabilities can be affected unevenly by pruning. To ensure targeted and balanced recovery, we need to identify and group data points that focus on similar capabilities. To achieve this goal, we hypothesize that distinct geometric topological structures of recovery instruction data in the high-dimensional semantic space may correspond to different aspects of LLM capabilities. This hypothesis is based on the intuition that instructions requiring similar capabilities are likely to cluster together in the semantic space, forming identifiable topological structures. In detail, we employ a two-step approach on the embedding space of instructions. First, we apply a diffusion kernel to SentenceBERT (Reimers & Gurevych, 2019) embeddings for manifold learning:
@@
Here, $e(x_{i})$ is the obtained low-dimensional manifold representation of instruction $x_{i}$. This process helps uncover the intrinsic geometric structure in the semantic space while reducing dimensionality and preserving non-linear relationships. Then, non-negative matrix factorization (NMF)-based spectral clustering (Ding et al., 2005) is conducted based on such $e(x_{i})$ to identify natural groupings of instructions that potentially correspond to different LLM capabilities, leading to $K$ clusters as follows:
@@
The detailed process of these two steps are provided as below. In the first step of manifold learning, we first obtain the SentenceBERT embedding of each instruction $x_{i}$. Then, an adjacency matrix $A$ is constructed based on the pairwise Euclidean distances of these embeddings: $A_{ij}=\exp(-\frac{\|\text{SentenceBERT}(x_{i})-\text{SentenceBERT}(x_{j})\|_{2}^{2}}{2\sigma^{2}})$, where $\sigma$ is a scaling parameter, typically set to the median of all pairwise distances. The degree matrix $D$ is then computed as a diagonal matrix where each diagonal element is the sum of the corresponding row in $A$:$D_{ii}=\sum_{j=1}^{n}A_{ij}$. Using these matrices, we define the normalized graph Laplacian $L=I-D^{-1/2}AD^{-1/2}$, where $I$ is the identity matrix. We then apply the diffusion kernel to this Laplacian $K_{t}=\exp(-tL)$, where $K_{t}$ is the diffusion kernel at time $t$. The diffusion time $t$ is selected using the spectral gap method: $t_{opt}=\operatorname*{arg\,max}_{t}\left(\frac{d\log(\lambda_{2}(t))}{d\log(t)}\right)$, where $\lambda_{2}(t)$ is the second eigenvalue of $K_{t}$. The low-dimensional manifold representation $e(x_{i})$ is then obtained by selecting the top $d$ eigenvectors of $K_{t_{opt}}$: $e(x_{i})=[\phi_{1}(x_{i}),\phi_{2}(x_{i}),...,\phi_{d}(x_{i})]$, where $\phi_{j}$ are the eigenvectors of $K_{t_{opt}}$ corresponding to the $d$ largest eigenvalues.
@@ 3.3 Capability Degradation-aware Instruction Selection
Capability Degradation Assessment To prioritize the severely affected capabilities and finally achieve the balanced recovery of pruned LLMs, we need a measure of capability degradation to guide the data selection. For each cluster $c_{k}$ obtained in Section 3.2, we define the capability degradation score (CDS) with the Jensen-Shannon divergence (JSD) (Fuglede & Topsoe, 2004) as follows:
$\frac{1}{|c_{k}|}\sum_{(x,y)\in c_{k}}\frac{1}{|y|}\sum_{m=1}^{|y|}\text{JSD}(P(t_{m}|M_{p},x)||P(t_{m}|M_{o},x)).$ (4)
Here, $P(t_{m}|M_{p},x)$ represents the output probability distribution on the $m$-th token of the pruned model $M_{p}$ given input $x$. Taking a token $t^{i}_{m}(1\leq i\leq|\text{Voc}|)$ in this distribution as an example, its corresponding probability is:
$P(t^{i}_{m}|M_{p},x)=\frac{\text{exp}(\frac{\text{logit}(t^{i}_{m})}{\tau})}{\sum^{|\text{Voc}|}_{j=1}\text{exp}(\frac{\text{logit}(t^{j}_{m})}{\tau})},$ (5)
@@
Thus, the obtained CDS quantifies the average performance degradation for data points in each capability cluster. The choice of JSD over simple loss variations as the performance degradation signal is motivated by its unique properties. First, its symmetry ensures consistent comparison between the pruned model $M_{p}$ and the original model $M_{o}$, while its bounded range (0 to 1) provides a normalized measure of divergence. This facilitates easier interpretation and comparison across different capability clusters. Moreover, JSD’s robustness to outliers and its information-theoretic foundation allow for a more nuanced assessment of capability degradation, capturing subtle changes in model outputs that might not be apparent from loss or accuracy values alone (Dutta et al., 2024) due to the sampling uncertainty. The smoothing effect introduced by the use of the mixed distribution in JSD calculation also contributes to a more stable assessment across diverse instruction types. By employing JSD in our CDS calculation, we obtain a comprehensive and reliable assessment of capability degradation, enabling more accurate identification and prioritization of the capabilities most severely affected by model pruning.
Inter-Capability Budget Allocation Sampling a subset of high-quality data from the original set to achieve better training performance is the objective of the data selection process. To ensure the efficiency on data utilization and training process, the instruction data budget $B(B<N)$ should be maintained. Under this budget, we design an adaptive selection approach based on the above CDS for balanced recovery while focusing on severely affected capabilities. In detail, we allocate sampling budget to each cluster proportionally to its corresponding CDS:
@@
Intra-Capability Efficiency-Driven Sample Selection To maximize computational efficiency during the recovery post-training phase, we need to select samples that offer the highest recovery benefit relative to their computational cost. Within each cluster $c_{k}$, we select top $n_{k}$ samples based on their Individual Efficiency Scores (IES):
$\text{IES}(x,y)=\frac{\frac{1}{|y|}\sum_{m=1}^{|y|}\text{JSD}(P(t_{m}|M_{p},x)||P(t_{m}|M_{o},x))}{\log\text{ComputationalCost}(x,y)}.$ (7)
@@
3.4 Negative Transfer Mitigation
To prevent performance degradation due to conflicting or irrelevant information, we need to detect and mitigate potential negative transfer. We introduce a Concept Consistency Graph (CCG) to model relationships between concepts in the selected data. Here, a concept refers to a key knowledge unit or semantic element extracted from an instruction tuning sample. Concepts play a crucial role in capturing the essential information within instructions and help to identify potential conflicts that could lead to negative transfer. By managing relationships between concepts, we aim to maintain semantic consistency across the selected instruction tuning dataset, thereby reducing the risk of learning conflicting or irrelevant information during the recovery process. The formal definition for CCG is provided as follows:

### v2

@@ Abstract
Model pruning is an effective approach for compressing large language models (LLMs). However, this process often leads to significant degradation of model capabilities. While post-training techniques such as instruction tuning are commonly employed to recover model performance, existing methods often overlook the uneven deterioration of model capabilities and incur high computational costs. Moreover, some irrelevant instructions may also introduce negative effects to model capacity recovery. To address these challenges, we propose the Post-training dAta Selection method for Efficient pruned large language model Recovery (PASER). PASER aims to identify instructions to recover the most compromised model capacities with a certain data budget. Our approach first applies manifold learning and spectral clustering to group recovery instructions in the semantic space, revealing capability-specific instruction sets. Then, the data budget is adaptively allocated across clusters by the degree of corresponding model capability degradation. In each cluster, we prioritize data samples that lead to the most decline of model performance. To mitigate potential negative tuning effects, we also detect and filter out conflicting or irrelevant recovery data. Extensive experiments demonstrate that PASER significantly outperforms conventional baselines, effectively recovering the general capabilities of pruned LLMs while utilizing merely 4%-20% of the original post-training data. We provide the anonymous code repository in Link.
@@ 1 Introduction
Model pruning, which aims at reducing model parameter amounts while maintaining model capabilities, has been a promising approach for large language model (LLM) compression. Mainstream LLM pruning schemes are unstructured (Frantar and Alistarh, 2023), semi-structured (Sun et al., 2024), and structured pruning (Ma et al., 2023). Even though effective, the model capability degradation after pruning is almost unavoidable, especially under high pruning ratios. This degradation phenomenon is often more severe for the structured pruning scheme (Dong et al., 2024), which has been widely adopted in industrial LLM compression thanks to its hardware-friendly property. Therefore, recovery post-training after pruning has been a silver bullet to address thisissue (Ma et al., 2023; Zhao et al., ). Among various types of data including pre-training corpora and extensive fine-tuning datasets (Xia et al., 2024; Sun et al., 2024), instruction tuning data has demonstrated unique advantages for efficient capability recovery (Ma et al., 2023; Zhao et al., ; Zhang et al., 2024; Chen et al., 2023). Compared to recovery via continued pre-training which requires massive computational resources, instruction tuning enables effective restoration with a much smaller data scale by explicit supervision. Furthermore, through the diverse task coverage, like language modeling, common sense reasoning, mathematical problem solving, and code generation (Taori et al., 2023; Wu et al., 2024), instruction tuning preserves the model’s general-purpose capabilities while preventing over-specialized recovery.
@@
Conventional schemes (Ma et al., 2023) usually employ the full instruction tuning datasets like Alpaca (Taori et al., 2023) to conduct the recovery post-training. However, this can bring significant computation overhead and even unsatisfactory recovery performance (See Appendix A). An intuitive solution is to leverage part of the original data for training, thus consuming less data and correspondingly reducing the computation resource demand. Nevertheless, directly utilizing the naively split data subset (e.g., first 20% of the data), can lead to sub-optimal performance, or even further performance degradation. Moreover, the recovered performance considerably varies for models trained with different subsets, implying the sensitivity to data constitution. Therefore, selecting the most valuable instruction-tuning data that can contribute to recovery performance positively and reduce training costs becomes crucial. Though previous works have noticed the significance of selecting high-quality data for general instruction tuning (Wang et al., 2024; Cao et al., ; Li et al., 2024a), few of them are specifically designed for the recovery post-training. Note such general high-quality standards, such as clear structure and natural, human-like expressions, may not effectively target the specific capabilities (e.g., math, code) severely compromised during pruning, which is essential for recovery post-training.
@@
Considering the above limitations, the ideal recovery post-training data selection approach should exhibit the following properties: (1) Targeted and Balanced Capability Recovery: Given the uneven deterioration of different capabilities in the pruning (see Appendix A), the ideal selection method should identify and prioritize the severely-degraded ones, while ensuring balanced recovery of the model’s overall functionality. (2) Recovery Training Efficiency: Limited computing resources pose serious efficiency challenges to the LLM recovery post-training. An ideal method should be able to select instructions that are both most beneficial for recovery and light in computational cost, thereby accelerating the training process and optimizing resource utilization. (3) Mitigation of Negative Tuning Effects: Recognizing that not all instruction data is beneficial for model recovery, an optimal approach should not only identify the most advantageous instructions, but also filter out potentially harmful or irrelevant ones. This significantly reduces the risk of negative tuning effects during the recovery training, ensuring that the selected data contributes positively to model recovery.
@@
To achieve such goals, we propose the Post-training dAta Selection method for Efficient pruned large language model Recovery (PASER). First, we perform semantic-structural recovery instruction clustering to obtain data groups corresponding to different LLM capabilities. Second, we select recovery instructions in a capability degradation-aware manner, where the data budget is allocated across different clusters based on their corresponding capability degradation degrees. In particular, when selecting samples within each capability cluster, the post-training computation cost of each sample is also taken into consideration to optimize the efficiency of the recovery process. Finally, we construct the concept consistency graph to maintain the semantic consistency across selected instructions, thus preventing the introduction of conflicting or irrelevant samples. We take the LLaMA 2/3/3.1, Baichuan2, Qwen2.5/3, and Mixtral 8$\times$7B as the target LLMs and perform the experiments under different LLM pruning schemes and different-sized instruction tuning datasets. The comparison with random and conventional instruction tuning data selection baselines demonstrates that PASER can more effectively enhance the recovered LLM performance on language modeling and various reasoning tasks. Meanwhile, the recovery training overhead can also be reduced significantly.
@@ 2 Related Works
Large Language Model Pruning can be generally divided into three categories: unstructured pruning, semi-structured pruning, and structured pruning. Unstructured pruning removes individual weights without structural constraints, with representative works including SparseGPT (Frantar and Alistarh, 2023), Wanda (Sun et al., 2024), BESA (Xu et al., 2024b), and OWL (Yin et al., 2024). This technique allows for maximum flexibility in weight selection and can achieve high compression rates while maintaining model performance. However, the resulting irregular sparsity patterns limits the practical acceleration. Semi-structured pruning (Guo et al., 2024; Malla et al., 2024; Frantar and Alistarh, 2023; Sun et al., 2024) targets specific patterns like N:M sparsity, balancing flexibility and hardware efficiency. Structured pruning approaches like LLM-Pruner (Ma et al., 2023) and SliceGPT (Ashkboos et al., 2024) remove entire structural components, offering better hardware compatibility and attracting industry attention (Ko et al., 2023; An et al., 2024; Song et al., 2024; Xia et al., 2024). However, structured pruning faces more severe performance degradation, highlighting the importance of recovery post-training.
@@ 3 Methodology
In this section, we first formulate the problem and then introduce three main components of the PASER framework (shown in Figure 1): semantic-structural recovery instruction clustering, capability degradation-aware instruction selection, and negative tuning effects mitigation. Furthermore, we provide the time complexity analysis for the PASER process.
Figure 1: Visualization for our proposed PASER recovery post-training data selection framework.
@@ 3.1 Problem Formulation
Let $M_{o}$ denote the original large language model and $M_{p}$ the pruned version of this model. We define the instruction tuning dataset as $D=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ represents an instruction and $y_{i}$ is its corresponding output. Our goal is to select a subset $S\subset D$ to efficiently recover the performance of $M_{p}$. We formulate the problem as an optimization task:
@@ 3.2 Semantic-Structural Recovery Instruction Clustering
After the LLM pruning, different model capabilities can be affected unevenly because such capabilities rely on distinct sets of internal modules (Meng et al., 2022; Dai et al., 2022), which are selectively removed. To ensure targeted and balanced recovery, we need to identify and group data points related to similar capabilities. To achieve this goal, we hypothesize that distinct geometric structures of recovery instruction data in the high-dimensional semantic space may correspond to different aspects of LLM capabilities. This hypothesis is based on the intuition that instructions pertaining similar capabilities are likely to cluster together in the semantic space, forming identifiable topological structures. Thus, we employ a two-step approach on the embedding space of instructions. First, to enhance the natural separation between capability-specific instruction clusters, we apply a diffusion kernel to SentenceBERT (Reimers and Gurevych, 2019) embeddings for manifold learning:
@@
Here, $e(x_{i})$ is the obtained low-dimensional manifold representation of instruction $x_{i}$. This process helps uncover the intrinsic geometric structure in the semantic space while reducing dimensionality and preserving non-linear relationships. Then, non-negative matrix factorization (NMF)-based spectral clustering (Ding et al., 2005) is conducted based on such $e(x_{i})$ to identify natural groupings of instructions that potentially correspond to different LLM capabilities, leading to $K$ clusters:
@@
The details are provided as below. In the first step of manifold learning, we first obtain the SentenceBERT embedding of each instruction $x_{i}$. Then, an adjacency matrix $A$ is constructed based on the pairwise Euclidean distances of these embeddings: $A_{ij}=\exp(-\frac{\|\text{SentenceBERT}(x_{i})-\text{SentenceBERT}(x_{j})\|_{2}^{2}}{2\sigma^{2}})$, where $\sigma$ is a scaling parameter, typically set to the median of all pairwise distances. The degree matrix $D$ is then computed as a diagonal matrix where each diagonal element is the sum of the corresponding row in $A$:$D_{ii}=\sum_{j=1}^{n}A_{ij}$. Using these matrices, we define the normalized graph Laplacian $L=I-D^{-1/2}AD^{-1/2}$, where $I$ is the identity matrix. We then apply the diffusion kernel to this Laplacian $K_{t}=\exp(-tL)$, where $K_{t}$ is the diffusion kernel at time $t$. The diffusion time $t$ is selected using the spectral gap method: $t_{opt}=\operatorname*{arg\,max}_{t}\left(\frac{d\log(\lambda_{2}(t))}{d\log(t)}\right)$, where $\lambda_{2}(t)$ is the second eigenvalue of $K_{t}$. The low-dimensional manifold representation $e(x_{i})$ is then obtained by selecting the top $d$ eigenvectors of $K_{t_{opt}}$: $e(x_{i})=[\phi_{1}(x_{i}),\phi_{2}(x_{i}),...,\phi_{d}(x_{i})]$, where $\phi_{j}$ are the eigenvectors of $K_{t_{opt}}$ corresponding to the $d$ largest eigenvalues.
@@ 3.3 Capability Degradation-Aware Instruction Selection
Capability Degradation Assessment To prioritize the severely affected capabilities and finally achieve the balanced recovery of pruned LLMs, we need a measure of capability degradation to guide the data selection. For each cluster $c_{k}$ obtained in Section 3.2, we define the capability degradation score (CDS) with the Jensen-Shannon divergence (JSD) (Fuglede and Topsoe, 2004), which measures distributional differences between the original and pruned models’ outputs as follows:
$\frac{1}{|c_{k}|}\sum_{(x,y)\in c_{k}}\frac{1}{|y|}\sum_{m=1}^{|y|}\text{JSD}(P(\mathbf{t}_{m}|M_{p};x,y_{<m})||P(\mathbf{t}_{m}|M_{o};x,y_{<m})).$ (4)
Here, $P(\mathbf{t}_{m}|M_{p};x,y_{<m})$ represents the output probability distribution on the $m$-th token variable of the pruned model $M_{p}$ given input $x$ and previous $m-1$ tokens of output $y$. Taking a token value $t^{i}_{m}(1\leq i\leq|\text{Voc}|)$ in this distribution as an example, its corresponding probability is
$P(t^{i}_{m}|M_{p};x,y_{<m})=\frac{\text{exp}(\frac{\text{logit}(t^{i}_{m})}{\tau})}{\sum^{|\text{Voc}|}_{j=1}\text{exp}(\frac{\text{logit}(t^{j}_{m})}{\tau})},$ (5)
@@
Inter-Capability Budget Allocation Sampling a subset of high-quality data $S$ from the original set $D$ to achieve better training performance is the objective of the data selection process. Typically, an instruction data budget $B(B<N)$ should be maintained to ensure the efficiency on data utilization and training process. Under this budget, we design an adaptive selection approach based on the above CDS for balanced recovery while focusing on severely affected capabilities. In detail, we allocate the sampling budget to each cluster proportionally to its corresponding CDS:
@@
Intra-Capability Efficiency-Driven Sample Selection To maximize computational efficiency during the recovery post-training phase, we need to select samples that offer the highest recovery benefit relative to their computational cost. Within each cluster $c_{k}$, we select the top $n_{k}$ samples based on their Individual Efficiency Scores (IES):
$\text{IES}(x,y)=\frac{\frac{1}{|y|}\sum_{m=1}^{|y|}\text{JSD}(P(\mathbf{t}_{m}|M_{p};x,y_{<m})||P(\mathbf{t}_{m}|M_{o};x,y_{<m}))}{\log\text{ComputationalCost}(x,y)}.$ (7)
@@
3.4 Negative Tuning Effects Mitigation
To prevent performance degradation due to conflicting or irrelevant information, we need to detect and mitigate the potential negative tuning effects. We introduce a Concept Consistency Graph (CCG) to model relationships between concepts in the selected data. Here, a concept refers to a key knowledge unit or semantic element extracted from an instruction tuning sample. Concepts play a crucial role in capturing the essential information within instructions and help to identify potential conflicts that could lead to a negative tuning effects. By managing relationships between concepts, we aim to maintain semantic consistency across the selected instruction tuning dataset, thereby reducing the risk of learning conflicting or irrelevant information during the recovery process. The formal definition for CCG is provided as follows: